OpenStudy (babyslapmafro):

How do you rationalize the following numerator? Click to see equation...

OpenStudy (babyslapmafro):

\[\frac{ \sqrt{x+4}-2 }{ x }\]

OpenStudy (anonymous):

multiply top and bottom by its conjugate

OpenStudy (anonymous):

\[\sqrt{x+4}+2\]

OpenStudy (anonymous):

can you finish from here?

OpenStudy (babyslapmafro):

I will try

OpenStudy (anonymous):

i like your answer